# Graph-theoretical estimates of the diameters of the Rubik’s Cube groups

### v2

@@ Abstract
A strict lower bound for the diameter of a symmetric graph is proposed, which is calculable with the order ($n$) and other local parameters of the graph such as the degree, even girth $g$ ($\geq 4$), and number of $g$-cycles traversing a vertex, which are easily determined by inspecting a small portion of the graph (unless the girth is large). It is applied to the symmetric Cayley graphs of the Rubik’s Cube groups of various sizes and metrics, yielding slightly tighter lower bounds than those of random $k$-regular graphs proposed by Bollobás and de la Vega, which range from 60% to 77% of the correct diameters of large-$n$ graphs.
@@ I Introduction
Figure 1: Generalized Petersen graph $G(4,1)$ as the Cayley graph of the $3\times 3\times 3$ Rubik’s Cube group in the square-slice-turn metric. Each vertex is a unique configuration (a pattern of face colors of the cubies while holding the orientation of the whole Cube fixed) reachable by a combination of the $R^{2}L^{-2}$, $D^{2}U^{-2}$, and $B^{2}F^{-2}$ turns (in the Singmaster notationSingmaster (1981)) from the solved configuration. Each edge connecting two vertexes is a turn (and its inverse) that brings one configuration denoted by one of the vertexes to another represented by the other vertex. The order ($n$) of the graph is 8, the girth ($g$) is 4, the diameter ($d$) is 3, and the number ($\eta$) of $g$-cycles traversing a vertex is 3.
Figure 2: Generalized Petersen graph $G(12,5)$ as the Cayley graph of the $2\times 2\times 2$ Rubik’s Cube group in the square-turn metric. The order ($n$) is 24, the girth ($g$) is 6, the diameter ($d$) is 4, and the number ($\eta$) of $g$-cycles per vertex is 3.
The fewest number of turns to solve the Rubik’s Cube from its hardest initial configuration is colloquially referred to as God’s number.Hofstadter (1981); Singmaster (1981); Joyner (2008); Rokicki et al. (2013, 2014a, 2014b) It is equal to the diameter of the Cayley graph of the Rubik’s Cube group. Previously,Hirata (2024) we estimated these diameters for the Cubes of various sizes in different metrics (a set of allowed combinations of rotations, each counted as one turn) using probabilistic logic. The estimates are accurate to within 2 of the actual values and are in exact agreement with the latter for both the 2$\times$2$\times$2 and 3$\times$3$\times$3 Cubes when the quarter-turn metric is adopted.
The Cayley graphs of the Rubik’s Cube group in some metrics are a symmetric graph.Biggs (1993); West (2019) For example, the 3$\times$3$\times$3 Cube group in the square-slice-turn metric ($R^{2}L^{-2}$, $D^{2}U^{-2}$, and $B^{2}F^{-2}$ in the Singmaster notationSingmaster (1981)) has the Cayley graph that is a cubic symmetric graph known as the generalized Petersen graph $G(4,1)$ shown in Fig. 1.West (2019) The Cayley graph of the 2$\times$2$\times$2 Cube group in the square-turn metricSingmaster (1981) ($R^{2}$, $D^{2}$, and $B^{2}$) is another cubic symmetric graph, which is identified as the generalized Petersen graph $G(12,5)$ (Fig. 2).West (2019)
@@
In this article, we introduce the formula for a strict lower bound for the diameter of a general symmetric graph derived with algebraic-graph-theoretical arguments.Biggs (1993); West (2019) It can be evaluated with the order and some easily determined local parameters of the graph, such as the degree, girth ($g$), and number of $g$-cycles traversing a vertex. We then apply this formula to the diameters of the Cube groups of different sizes and metrics and compare the lower bounds with the probabilistic estimatesHirata (2024) and actual values. See Ref. Hirata, 2024 for nomenclature of the Rubik’s Cube groups.
@@ II Symmetric Cayley graphs
Following the notation of Biggs,Biggs (1993) a distance-partition graph $\Gamma_{i}(v)$ for any vertex $v$ of a symmetric graph $\Gamma$ of order $n$ and even girth $g\,(\geq 4)$ is defined by its vertex set,
@@
at any $i$ ($0\leq i\leq d$). Its corresponding edge set is defined by
@@
at any $i$. If the graph $\Gamma$ has a diameter $d$,
$|E\Gamma_{d}(v)|=0,$ (5)
@@
The $d_{i}$ ($0\leq i\leq d$) is therefore the number of vertexes with distance $i$ from an arbitrary origin vertex; $d_{0}=1$ and $d_{1}=k$. Let $\bm{r}$ be an array of “branching ratios” defined by
@@
The $r_{i}$ ($1\leq i\leq d$) is the growth rate of the number of vertexes with increasing distance, i.e., $r_{1}=k$, $r_{2}=k-1$, and generally,
@@
These can be understood as follows: From the origin vertex (distance 0), $k$ edges emanate and reach $k$ new vertexes at distance 1 ($r_{1}=k$). From each of the $k$ vertexes at distance 1, one edge connects the origin vertex and the remaining $k-1$ edges spawn $k-1$ new vertexes at distance 2 ($r_{2}=k-1$). From each of the $k-1$ vertexes at distance 2, at most $k-1$ new vertexes at distance 3 can be reached ($r_{3}\leq k-1$). The equal sign applies when the girth is greater than 4. When the girth is equal to 4, some of the vertexes are a shared destination of two edges and therefore the number of new vertexes at distance 3 is less ($r_{3}<k-1$). In subsequent steps, possible existence of longer cycles decreases the branching ratio even further ($r_{i+1}\leq r_{i}$).
@@ III Diameters of symmetric graphs
In a symmetric graph with order $n$ and degree $k$, let $g$ ($\geq 4$) be the even girth, $d$ be the diameter, and $\eta$ be the number of cycles of length $g$ traversing a vertex. We seek a lower bound for $d$, whose precise determination is generally hard, given in terms of $n$ and local parameters of the graph—$k$, $g$, and $\eta$—which are usually readily available.
@@
$n_{\text{max}}$ $=$ $\left\lfloor n_{0}+k(k-1)^{g/2-2}\frac{r_{\text{max}}^{d-g/2+2}-r_{\text{max}}}{r_{\text{max}}-1}\right\rfloor,$ (16)
@@
which is related to the Moore bound (and equal to the latter when $g=2d$ and $\eta=0$).Biggs (1993); West (2019) When $r_{\text{max}}=1$, we instead have
@@
We have confirmed that they indeed give a lower bound for the diameter for each of the cubic symmetric graphs (with an even girth, conservatively assuming $\eta=1$) listed in a Foster Census.Conder and Dobcsányi (2002)
@@
In our previous study,Hirata (2024) the diameter of a Rubik’s Cube group is estimated by the following formula,
$d_{\text{probab}}$ $=$ $\frac{\ln n}{\ln r_{\text{max}}}+\frac{\ln n}{r_{\text{max}}}.$ (21)
which is neither a lower nor upper bound, but works for nonsymmetric Cayley graphs or for any metric that does not form a mathematically well-defined group. It has been derived on the basis of a probabilistic argument or a modified version of the coupon collector’s problem.Hirata (2024) Equation (21) can also be viewed as a probabilistic estimate of the diameter of a symmetric group, where $r_{\text{max}}$ is given by Eq. (14) (in Ref. Hirata, 2024, a more accurate estimate of $r$ was computationally determined and used instead). The first term approximately counts the distance of the growth phase, where $r\approx r_{\text{max}}$ and the number of vertexes increases geometrically. The second term measures the distance of the subsequent rapid decay phase. Our graph-theoretical lower bound [Eq. (LABEL:dmin1)] considers only the growth phase and its formula thus has a similar functional form with the first term of Eq. (21); they become nearly identical in the $n\to\infty$ limit.
Upper and lower bounds for the diameter of random $k$-regular graphs were obtained by Bollobás and de la Vega.Bollobás and de la Vega (1982); Bollobás (2001) For “almost every” graph and a constant $\epsilon>0$, the BV bounds are
$d_{\text{min}}^{\text{BV}}$ $=$ $\left\lfloor\log_{k-1}n\right\rfloor+\left\lfloor\log_{k-1}\ln n-\log_{k-1}\frac{6k}{k-2}\right\rfloor+1,$ (22)
$d_{\text{max}}^{\text{BV}}$ $=$ $\left\lceil{\log_{k-1}\left\{(2+\epsilon)\,kn\ln n\right\}}\right\rceil+1,$ (23)
@@ IV Cayley graphs of Rubik’s Cube groups
The reader is referred to our previous reportHirata (2024) for the terminologies, group/graph orders, and other details of the Rubik’s Cube groups of various sizes and metrics. Table 1 summarizes the results of applying above formulas.
@@
Cube Metric Order $n$ Degree $k$ Girth $g$ $\eta$ [^1] The number of $g$-cycles per vertex. $r_{\text{max}}$ [^2] Upper bound for the branching ratio at distance $g/2-1$ and greater [Eq. (14)]. $d$ (actual) [^3] The diameter of the 3$\times$3$\times$3 Cube group in the quarter-turn metric was determined computationally by Rokicki et al.Rokicki et al. (2014b, 2013, a) Other diameters listed are also determined computationally more easily. $d_{\text{min}}$ [^4] Graph-theoretical lower bound for the diameter $d$ [Eq. (LABEL:dmin1) or (20)] (this work). $d_{\text{probab}}$ [^5] Probabilistic estimate [Eq. (21)].Hirata (2024) $d_{\text{min}}^{\text{BV}}$ [^6] Asymptotic ($n\to\infty$) lower bound [Eq. (22)] and upper bound ($\epsilon=0$) [Eq. (23)] of Bollobás and de la Vega.Bollobás and de la Vega (1982); Bollobás (2001) $d_{\text{max}}^{\text{BV}}$ [^6] Asymptotic ($n\to\infty$) lower bound [Eq. (22)] and upper bound ($\epsilon=0$) [Eq. (23)] of Bollobás and de la Vega.Bollobás and de la Vega (1982); Bollobás (2001)
@@ IV.1 The 3$\times$3$\times$3 Cube in the square-slice-turn metric
The order of the graph $n$ is a global parameter and is potentially hard to determine. However, for a Rubik’s Cube group, it can be worked out by a careful combinatorial argumentHofstadter (1981); Singmaster (1981); Hirata (2024) or with the aid of group theory.Joyner (2008) In the square-slice metric,Singmaster (1981) each of the three center slices is allowed to independently take one of two orientations and thus $n=2^{3}=8$. Other local parameters such as the degree $k$, girth $g$, and the number $\eta$ of $g$-cycles per vertex can be gleaned by inspecting a small neighborhood of a vertex in the partial graph. From Fig. 1, we find that $k=3$, $g=4$, and $\eta=3$.
@@
Using Eq. (14), $r_{\text{max}}=1$, which corresponds to the predicted distance array of $\bm{d}=\{1,3,3,3\}$. Equation (20) yields $d_{\text{min}}=\lceil 2.3\rceil=3$, which is a tight lower bound for the actual diameter of 3. The probabilistic estimate [Eq. (21)] breaks down in this tiny case because $n=8$ is too few for probabilistic logic. For the same reason, the BV lower and upper bounds [Eqs. (22) and (23)] are too low and too high to be useful.
@@ IV.3 The 2$\times$2$\times$2 Cube in the quarter-turn metric
Figure 3: A partial Cayley graph of the 2$\times$2$\times$2 Rubik’s Cube group in the quarter-turn metric, which is symmetric with degree ($k$) 6. The order ($n$) is $3.67\times 10^{6}$, the girth ($g$) is 4, the diameter ($d$) is 14, and the number ($\eta$) of $g$-cycles per vertex is 3. The vertexes on the periphery are further connected to other vertexes that are not shown.
Figure 3 shows a local Cayley graph within distance 3 from an arbitrary origin vertex for the 2$\times$2$\times$2 Cube group in the quarter-turn metric.Singmaster (1981) The order of the graph can be determined by combinatorial logic to be $n=3.67\times 10^{6}$.Hofstadter (1981); Singmaster (1981); Joyner (2008); Hirata (2024)
Without drawing such a huge graph, based on the facts that all Cayley graphs are vertex-transitive and that all turns are symmetrically equivalent with one another, the Cayley graph of this group should also be edge-transitive and thus symmetric.Biggs (1993); West (2019) This is supported by the appearance of a partial graph (within distance 3 from an arbitrarily chosen origin vertex) drawn in Fig. 3, in which every vertex is equivalent and interchangeable with one another, and so is every edge. This smallest portion of the Cayley graph is sufficient to determine $k=6$, $g=4$, and $\eta=3$.
@@
Unlike these easily determined parameters, the diameter ($d$) is hard to come by. For the 2$\times$2$\times$2 Cube, a brute-force computation of the distance array $\bm{d}$ is possible,Hirata (2024) as shown in Fig. 4. The computed $\bm{d}$ indicates that Eq. (8) is obeyed and $d=14$. It is our objective, however, to estimate $d$ from the easily obtained local information of the graph as well as $n$.
Using Eq. (LABEL:dmin1), we find $d_{\text{min}}=\lceil 9.7\rceil=10$, which bounds the actual $d=14$ from below as it should. The probabilistic estimate [Eq. (21)] is $d_{\text{probab}}=13.4$, which is much closer to the correct value of 14, as expected. The BV lower bound $d_{\text{min}}^{\text{BV}}=10$ [Eq. (22)] is the same as $d_{\text{min}}$, which is not surprising considering the similarity in their underlying logic. The $d_{\text{max}}^{\text{BV}}=14$ is a tight upper bound and equal to the actual diameter ($d=14$) itself, but this is likely the result of error cancellation (see below). On the other hand, the good agreement between $d_{\text{probab}}=13.4$ and the actual $d$ is not accidental.Hirata (2024)

### v3

@@ Abstract
A strict lower bound for the diameter of a symmetric graph is proposed, which is calculable with the order ($n$) and other local parameters of the graph such as the degree ($k$), even girth $g$ ($\geq 4$), and number of $g$-cycles traversing a vertex, which are easily determined by inspecting a small portion of the graph (unless the girth is large). It is applied to the symmetric Cayley graphs of some Rubik’s Cube groups of various sizes and metrics, yielding slightly tighter lower bounds of the diameters than those for random $k$-regular graphs proposed by Bollobás and de la Vega. They range from 60% to 77% of the correct diameters of large-$n$ graphs.
@@ I Introduction
Figure 1: Generalized Petersen graph $G(4,1)$ as the Cayley graph of the $3\times 3\times 3$ Rubik’s Cube group in the square-slice-turn metric. Each vertex is a unique configuration (a pattern of face colors of the cubies while holding the orientation of the whole Cube fixed) reachable by a combination of the $R^{2}L^{-2}$, $D^{2}U^{-2}$, and $B^{2}F^{-2}$ turns (in the Singmaster notation Singmaster (1981)) from the solved configuration. Each edge connecting two vertexes is a turn (and its inverse) that brings one configuration denoted by one of the vertexes to another represented by the other vertex. The order of the graph is 8, the degree is 3, the girth ($g$) is 4, the diameter is 3, and the number of $g$-cycles traversing a vertex is 3.
Figure 2: Generalized Petersen graph $G(12,5)$ as the Cayley graph of the $2\times 2\times 2$ Rubik’s Cube group in the square-turn metric. The order is 24, the degree is 3, the girth ($g$) is 6, the diameter is 4, and the number of $g$-cycles per vertex is 3.
The fewest number of turns to solve the Rubik’s Cube from its hardest initial configuration is colloquially referred to as God’s number Hofstadter (1981); Singmaster (1981); Joyner (2008); Rokicki et al. (2013, 2014, 2024). It is equal to the diameter of the Cayley graph of the Rubik’s Cube group. In the preceding article Hirata (2024a), we estimated these diameters for the Cubes of various sizes in different metrics (a set of allowed combinations of rotations, each counted as one turn) using probabilistic logic. The estimates are accurate to within 2 of the actual values and are in exact agreement with the latter for both the 2$\times$2$\times$2 and 3$\times$3$\times$3 Cubes when the quarter-turn metric is adopted.
The Cayley graphs of the Rubik’s Cube group in some metrics are a symmetric graph Bollobás (1978); Biggs (1993); Bollobás (2001); West (2019). For example, the 3$\times$3$\times$3 Cube group in the square-slice-turn metric ($R^{2}L^{-2}$, $D^{2}U^{-2}$, and $B^{2}F^{-2}$ in the Singmaster notation Singmaster (1981)) has the Cayley graph that is a cubic symmetric graph known as the generalized Petersen graph $G(4,1)$ shown in Fig. 1 West (2019). The Cayley graph of the 2$\times$2$\times$2 Cube group in the square-turn metric Singmaster (1981) ($R^{2}$, $D^{2}$, and $B^{2}$) is another cubic symmetric graph, which is identified as the generalized Petersen graph $G(12,5)$ (Fig. 2) West (2019).
@@
In this article, we introduce the formula for a strict lower bound for the diameter of a general symmetric graph derived with algebraic-graph-theoretical arguments Biggs (1993); West (2019). It can be evaluated with the order and some easily determined local parameters of the graph, such as the degree ($k$), girth ($g$), and number of $g$-cycles traversing a vertex. We then apply this formula to the diameters of the Cube groups of different sizes and metrics and compare the lower bounds with the probabilistic estimates Hirata (2024a) and actual values as well as with the lower and upper bounds of the diameter of a random $k$-regular graph proposed by Bollobás and de la Vega Bollobás and de la Vega (1982); Bollobás (2001). See the preceding article Hirata (2024a) for previous studies and nomenclature of the Rubik’s Cube groups.
@@ II Symmetric Cayley graphs
Following the notation of Biggs Biggs (1993), a distance-partition graph $\Gamma_{i}(v)$ for any vertex $v$ of a symmetric graph $\Gamma$ of order $n$, degree $k\,(\geq 3)$, and even girth $g\,(\geq 4)$ is defined by its vertex set,
@@
at any $i$ ($0\leq i\leq d$) because $\Gamma$ is a symmetric graph. Its corresponding edge set is defined by
@@
at any $i$. If $\Gamma$ has a diameter $d$,
$|E\Gamma_{d+1}(v)|=0,$ (5)
@@
The $d_{i}$ ($0\leq i\leq d$) is therefore the number of vertexes with distance $i$ from an arbitrary origin vertex; $d_{0}=1$ and $d_{1}=k$. Let $\bm{r}$ be an array of branching ratios defined by
@@
The $r_{i}$ ($1\leq i\leq d$) is the rate of growth in the number of vertexes with increasing distance, i.e., $r_{1}=k$, $r_{2}=k-1$, and generally,
@@
These can be understood as follows: From the origin vertex (distance 0), $k$ edges emanate and reach $k$ new vertexes at distance 1 ($r_{1}=k$). From each of the $k$ vertexes at distance 1, one edge connects the origin vertex, but the remaining $k-1$ edges spawn $k-1$ new vertexes at distance 2 ($r_{2}=k-1$). From each of the $k-1$ vertexes at distance 2, at most $k-1$ new vertexes at distance 3 can be reached ($r_{3}\leq k-1$). The equal sign applies when the girth is greater than 4. When the girth is equal to 4, some of the vertexes are a shared destination of two edges and therefore the number of new vertexes at distance 3 is less ($r_{3}<k-1$). In subsequent steps, possible existence of longer cycles decreases the branching ratio even further ($r_{i+1}\leq r_{i}$).
@@ III Diameters of symmetric graphs
In a symmetric graph with order $n$ and degree $k\,(\geq 3$), let $g\,(\geq 4$) be the even girth, $d$ be the diameter, and $\eta$ be the number of cycles of length $g$ traversing a vertex. We seek a lower bound for $d$, whose precise determination is generally hard, in terms of $n$ and local parameters of the graph—$k$, $g$, and $\eta$—which are usually readily available.
@@
$n_{\text{max}}$ $=$ $\left\lfloor n_{0}+k(k-1)^{g/2-2}\frac{r_{\text{max}}^{d-g/2+2}-r_{\text{max}}}{r_{\text{max}}-1}\right\rfloor$ (16)
@@
which is related to the Moore bound (and equal to the latter when $g=2d$ and $\eta=0$) Biggs (1993); Bollobás (2001); West (2019). When $r_{\text{max}}=1$, we instead have
@@
We have confirmed that they indeed give a lower bound for the diameter for each of the cubic symmetric graphs (with an even girth, conservatively assuming $\eta=1$) listed in a Foster Census Conder and Dobcsányi (2002).
@@
In our previous study Hirata (2024a), the diameter of a Rubik’s Cube group is estimated by the following formula,
$d_{\text{probab}}$ $=$ $\log_{r_{\text{max}}}n+\frac{\log n}{r_{\text{max}}}.$ (21)
which is neither a lower nor upper bound, but works for nonsymmetric Cayley graphs or for any turn metric that does not even form a mathematically well-defined group. It has been derived on the basis of a probabilistic argument or a modified version of the coupon collector’s problem Hirata (2024a). Equation (21) can also be viewed as a probabilistic estimate of the diameter of a symmetric group, where $r_{\text{max}}$ is given by Eq. (14) (in Ref. Hirata (2024a), a more accurate estimate of $r$ was computationally determined and used instead). The first term approximately counts the distance of the growth phase, where $r\approx r_{\text{max}}$ and the number of vertexes increases geometrically. The second term measures the distance of the subsequent rapid decay phase. Our graph-theoretical lower bound [Eq. (LABEL:dmin1)] considers only the growth phase and its formula thus has a similar functional form with the first term of Eq. (21); they become nearly identical in the $n\to\infty$ limit.
Upper and lower bounds for the diameter of a random $k$-regular graph were obtained by Bollobás and de la Vega Bollobás and de la Vega (1982); Bollobás (2001). For almost every graph and a constant $\epsilon>0$, the BV bounds are
$d_{\text{min}}^{\text{BV}}$ $=$ $\left\lfloor\log_{k-1}n\right\rfloor+\left\lfloor\log_{k-1}\log n-\log_{k-1}\frac{6k}{k-2}\right\rfloor+1,$ (22)
$d_{\text{max}}^{\text{BV}}$ $=$ $\left\lceil{\log_{k-1}\left\{(2+\epsilon)\,kn\log n\right\}}\right\rceil+1,$ (23)
@@ IV Cayley graphs of Rubik’s Cube groups
The reader is referred to the preceding article Hirata (2024a) for the terminologies, group/graph orders, and other details of the Rubik’s Cube groups of various sizes and metrics. Table 1 summarizes the results of applying above formulas.
@@
Cube Metric Order $n$ Degree $k$ Girth $g$ $\eta$ [^1] The number of $g$-cycles per vertex. $r_{\text{max}}$ [^2] Upper bound for the branching ratio at distance $g/2-1$ and greater [Eq. (14)]. $d$ (actual) [^3] The diameter of the 3$\times$3$\times$3 Cube group in the quarter-turn metric was determined computationally by Rokicki et al. Rokicki et al. (2024, 2013, 2014). Other diameters listed are easily determined computationally. $d_{\text{min}}$ [^4] Graph-theoretical lower bound for the diameter $d$ [Eq. (LABEL:dmin1) or (20)] (this work). $d_{\text{probab}}$ [^5] Probabilistic estimate [Eq. (21)] (the preceding article Hirata (2024a)). $d_{\text{min}}^{\text{BV}}$ [^6] Asymptotic ($n\to\infty$) lower bound [Eq. (22)] and upper bound ($\epsilon=0$) [Eq. (23)] of Bollobás and de la Vega Bollobás and de la Vega (1982); Bollobás (2001). $d_{\text{max}}^{\text{BV}}$ [^6] Asymptotic ($n\to\infty$) lower bound [Eq. (22)] and upper bound ($\epsilon=0$) [Eq. (23)] of Bollobás and de la Vega Bollobás and de la Vega (1982); Bollobás (2001).
@@ IV.1 The 3$\times$3$\times$3 Cube in the square-slice-turn metric
The order of the graph $n$ is a global parameter and is potentially hard to determine. However, for a Rubik’s Cube group, it can be worked out by a careful combinatorial argument Hofstadter (1981); Singmaster (1981); Hirata (2024a) or with the aid of group theory Joyner (2008). In the square-slice metric Singmaster (1981), each of the three center slices is allowed to independently take one of two orientations and thus $n=2^{3}=8$. Other local parameters such as the degree $k$, girth $g$, and the number ($\eta$) of $g$-cycles per vertex can be gleaned by inspecting a small neighborhood of a vertex in the partial graph. From Fig. 1, we find $k=3$, $g=4$, and $\eta=3$.
@@
Using Eq. (14), $r_{\text{max}}=1$, which corresponds to the predicted distance array of $\bm{d}=\{1,3,3,3\}$. Equation (20) yields $d_{\text{min}}=\lceil 2.3\rceil=3$, which is a tight lower bound for the actual diameter of 3. The probabilistic estimate [Eq. (21)] breaks down in this tiny case because $n=8$ is too few for probabilistic logic. For the same reason, the BV lower and upper bounds [Eqs. (22) and (23)] are too low and too high.
@@ IV.3 The 2$\times$2$\times$2 Cube in the quarter-turn metric
Figure 3: A partial Cayley graph of the 2$\times$2$\times$2 Rubik’s Cube group in the quarter-turn metric, which is symmetric with degree 6. The order is $3.67\times 10^{6}$, the girth ($g$) is 4, the diameter is 14, and the number of $g$-cycles per vertex is 3. The vertexes on the periphery are further connected to other vertexes that are not shown.
Figure 3 shows a local Cayley graph within distance 3 from an arbitrary origin vertex for the 2$\times$2$\times$2 Cube group in the quarter-turn metric Singmaster (1981). The order of the graph can be determined by combinatorial logic to be $n=3.67\times 10^{6}$ Hofstadter (1981); Singmaster (1981); Joyner (2008); Hirata (2024a).
Without drawing such a huge graph, based on the facts that all Cayley graphs are vertex-transitive and that all turns are symmetrically equivalent with one another, the Cayley graph of this group should also be edge-transitive and thus symmetric Biggs (1993); West (2019). This is supported by the appearance of a partial graph (within distance 3 from an arbitrarily chosen origin vertex) drawn in Fig. 3, in which every vertex is equivalent and interchangeable with one another, and so is every edge. This smallest portion of the Cayley graph is sufficient to determine $k=6$, $g=4$, and $\eta=3$.
@@
Unlike these easily determined parameters, the diameter ($d$) is hard to come by. For the 2$\times$2$\times$2 Cube, a brute-force computation of the distance array $\bm{d}$ is possible Hirata (2024a, b), as shown in Fig. 4. The computed $\bm{d}$ indicates that Eq. (8) is obeyed and $d=14$. It is our objective, however, to estimate $d$ from the easily obtained local information of the graph as well as $n$.
Using Eq. (LABEL:dmin1), we find $d_{\text{min}}=\lceil 9.7\rceil=10$, which bounds the actual $d=14$ from below as it should. The probabilistic estimate [Eq. (21)] is $d_{\text{probab}}=13.4$, which is much closer to the correct value of 14, as expected. The BV lower bound $d_{\text{min}}^{\text{BV}}=10$ [Eq. (22)] is the same as $d_{\text{min}}$, which is not surprising considering the similarity in their underlying derivations. The $d_{\text{max}}^{\text{BV}}=14$ is a tight upper bound and equal to the actual diameter ($d=14$) itself, but this is likely the result of error cancellation (see below). On the other hand, the good agreement between $d_{\text{probab}}=13.4$ and the actual $d$ is not accidental Hirata (2024a).